# Enhancing Feature Diversity Boosts Channel-Adaptive Vision Transformers

### v1

@@ 3.3 Diverse Channel Sampling (DCS)
Bao et al. introduced HCS to reduce the training time and improve the robustness of the model [18]. The main concept is to randomly drop some input channels and train the model only on the remaining channels. In the same spirit, we propose a novel method, Diverse Channel Sampling (DCS), to sample a more diverse subset of channels during training (Fig. 2(c)). Similar to HCS, we start by randomly sampling a number $n$, which is the size of a subset of channels to train on. However, while HCS samples $n$ channels randomly, DCS first samples an anchor channel $c_{n}$. Then, we select other $n-1$ channels that are dissimilar to the anchor channel. This idea shares similarity with Channel DropBlock [53], where a set of similar channels in a CNN layer is masked out to disrupt co-adapted features. However, instead of keeping a fixed number of feature map channels as in Channel DropBlock, DCS selects a flexible number of input channels for each sampling. The procedure of DCS is outlined in Algorithm 1.
@@
Temperature $t$
1 Sample a random variable $n$ uniformly from the set $\{1,2,...,m\}$
2 Sample an anchor channel $c_{n}$ uniformly from all $m$ channels
3 Compute the cosine similarity between channel $c_{n}$ and the other $m-1$ channels: $\mathbf{s}=[\left\langle f_{n},f_{i}\right\rangle,...]$, $\forall i\neq n$
4 Convert $1-\mathbf{s}$ to probability using $\mathrm{softmax}$ with temperature $t$: $\mathbf{p}=\mathrm{softmax}((1-\mathbf{s})/t)$
5 Sample $n-1$ distinct channels from $m-1$ channels with probability $\mathbf{p}$
6 Combine the $n-1$ channels with channel $c_{n}$ to create a set of $n$ sampled channels.
Output :  Image $X$ with only $n$ sampled channels
@@ 5 Sample $n-1$ distinct channels from $m-1$ channels with probability $\mathbf{p}$
In practice, Algorithm 1 can be applied to a batch of images for faster sampling. We use channel tokens to represent the channel feature $f$. The temperature $t$ controls the sharpness of the probability distribution. With a large $t$, DCS reduces to HCS, while with a small $t$, DCS selects a random subset of channels that are the least similar to the anchor channel.
@@ 3.4 Training Objective
The final loss is the combination of the main loss that is specific to the task (e.g., cross-entropy for classification), Channel Diversification Loss (CDL) applied to channel tokens, and Token Diversification Loss (TDL) used on patch tokens:
@@ 4.1 Experimental Setup
  • DepthwiseViT [14] uses a depthwise convolution layer, where each channel is filtered separately. The responses are averaged to create a feature representation, which is then fed into a ViT backbone.
@@
  • HyperNetViT [42] utilizes a Neural Network (e.g., MLP) to generate weights for each channel independently and then concatenate them to form a patch project layer, followed by a ViT backbone.
  • ChAda-ViT [15] uses a shared project layer to extract features from each channel separately, then feeds these tokens, together with their corresponding positional embeddings and channel embeddings, into a ViT backbone.
@@
Implementation details. As HCS proves robust in multi-channel imaging [18], we incorporate this technique for DepthwiseViT, TemplateMixingViT, and HyperNetViT to ensure a fair comparison in these adaptive baselines used by Chen et al. [14] [^1] https://github.com/chaudatascience/channel_adaptive_models. For ChannelViT and ChAda-ViT, due to their similarity (primarily a difference in whether HCS is included), we use the implementation from [18] for both methods [^2] https://github.com/insitro/ChannelViT. All baselines utilize a ViT small architecture ($21$M parameters) implemented in DINOv2 [54] as the backbone [^3] https://github.com/facebookresearch/dinov2. We use the AdamW optimizer [55] to train the models, minimizing cross-entropy loss on JUMP-CP and So2Sat, and proxy loss on CHAMMI. For the learning rate, we use a scheduler with linear warmup and cosine decay. Refer to Appendix Sec. A for details.
@@ 4.2 Datasets
So2Sat [17] contains synthetic aperture radar and multispectral optical image patches from remote sensing satellites. Each image in the dataset has $18$ channels, of which $8$ Sentinel-1 and $10$ Sentinel-2 channels. The dataset consists of $17$ classes, each representing a distinct climate zone. We use the city-split version of the dataset, which includes $352$K training images and $24$K test images.
@@ 4.3 Results
Table 1: Comparison of test accuracy of channel adaptive models. Our model outperforms other baselines, with a $5.0$% boost on CHAMMI and a $1.5-2.5$% point improvement on JUMP-CP and So2Sat. "Full" refers to inference on all channels, while "Partial" means testing on a subset of channels (Sentinel-1 channels for So2Sat, fluorescence channels for JUMP-CP).
@@
DiChaViT (ours) 69.68 69.24 57.98 63.36 47.76
Table 2: Test the accuracy of DiChaViT and ChannelViT on partial channels of JUMP-CP [12]. DiChaViT consistently exhibits improved robustness in the presence of missing channels during inference. The result on partial channels (mean±$\pm$±std) represents the average of all combinations.
@@
ChannelViT [18] 67.51 60.18$\pm$9.3 52.74$\pm$12.2 44.89$\pm$13.2 36.88$\pm$12.3 29.36$\pm$9.3 23.70$\pm$5.0 20.78$\pm$1.6
DiChaViT (ours) 69.24 62.19$\pm$9.2 54.49$\pm$12.4 46.35$\pm$13.4 38.00$\pm$12.4 30.09$\pm$9.3 23.97$\pm$4.9 20.90$\pm$1.6
Tab. 1 shows that DiChaViT outperforms the state-of-the-art ChannelViT by up to $5.0$% on all three datasets: CHAMMI [14], JUMP-CP [12], and So2Sat [17]. For JUMP-CP and So2Sat, we consider two scenarios: tested on all training channels (denoted as "Full") and tested on a subset of channels (denoted as "Partial"). In the full channels settings, our model shows a $1.5-2.5$% point improvement compared with other baselines on JUMP-CP and So2Sat. When tested on partial channels, DiChaViT demonstrates its robustness by achieving a $1.5$% improvement compared with the baselines. This demonstrates that diversifying feature representations in MCI-ViT models boosts both performance and robustness.
Tab. 2 presents a detailed evaluation of DiChaViT and the best baseline model, ChannelViT, when tested on partial channels of the JUMP-CP dataset (with a total of 8 channels). For the partial channel evaluation, we exclude some of the channels that the models were trained on and only test the model on the remaining channels. Then, we calculate the average accuracy across all combinations, e.g., testing on 7 channels involves averaging the results of $C^{7}_{8}=8$ combinations. We can observe that DiChaViT consistently demonstrates improved robustness when some input channels are missing.
4.4 Analysis and Discussion
The role of channel tokens. In MCI-ViT models such as ChannelViT [18] and ChAda-ViT [15], channel tokens play a crucial role in learning channel-specific features, particularly when dealing with multiple channels where each contains unique information. To assess the impact of channel tokens, we compared the performance of ChannelViT on JUMP-CP and CHAMMI with (orange bars) and without channel tokens (blue bars), as shown in Fig. 3. The results indicate that ChannelViT demonstrates significant improvements with channel tokens, resulting in $3.93$% and $14.88$% point increases on JUMP-CP and CHAMMI, respectively, highlighting their importance.
Figure 3: Performance of ChannelViT [18] on JUMP-CP and CHAMMI with and without channel tokens. Using channel tokens with orthogonal initialization (green) improves performance.
Orthogonal initialization of channel tokens boosts performance. As shown in Fig. 3, using orthogonal initialization (green) provides a $0.5-1$% gain on JUMP-CP and CHAMMI, indicating that the model benefits from more distinct channel feature representations.
Ablation study on DiChaViT’s architecture. Tab. 3 shows the performance of DiChaViT when removing each of its components. We notice that removing DCS has the most detrimental effect on performance, while CDL and TDL have a similar impact across the three datasets. The best performance is achieved when all components are used, indicating that each component plays a crucial role in the model’s design.
Table 3: Model ablations of DiChaViT. We evaluate the impact of each component in DiChaViT by removing it from the model. Removing any component has a negative impact on overall performance, with a significant decrease observed when DCS is removed. Including all components improves performance across all three datasets.
@@ 4.4 Analysis and Discussion
DiChaViT 69.68 69.24 57.98 63.36 47.76
w/o CDL 68.07 68.35 56.87 62.29 45.74
w/o TDL 68.25 68.00 56.62 62.39 46.87
w/o DCS 66.48 66.94 43.12 62.26 45.79
Table 4: Ablation on the two components of TDL. Only $\mathcal{L}_{\text{s}}$ indicates using only within channel tokens, while Only $\mathcal{L}_{\text{d}}$ indicates the use of only tokens from different channels in Eq. 4. Incorporating both components in TDL gives the best performance.
@@
Both 63.36 69.68
Table 5: Effect of temperature $t$ (Algorithm 1) on DCS. The first column ($\approx 0$) indicates the use of a very small value of $t$, which is reduced to selecting the lowest similarity channels. The last column indicates a large value of $t$, which is reduced to HCS [18]. Using $t=0.1$ obtain the best results on So2Sat and CHAMMI datasets.
Temperature $t$ $\approx 0$ 0.001 0.01 0.1 0.2 HCS
@@
CHAMMI [14] 67.22 66.91 68.96 69.68 66.07 66.30
Impact of $\lambda_{\mathrm{CDL}}$ (Eq. 5) in CDL. Fig. 4(a) and (b) show the performance of DiChaViT (mean and std) across different values of $\lambda_{\mathrm{CDL}}$ on So2Sat and CHAMMI. We can observe that selecting a value that is too large is not beneficial to the performance. It is worth finding a suitable value for $\lambda_{\mathrm{CDL}}$. On So2Sat, $\lambda_{\mathrm{CDL}}=0.001$ obtains the best performance, while the suitable value on CHAMMI is $0.1$.
Ablation on TDL (Eq. 4). Fig. 4(c) reports the performance of our model across different ratios of $\lambda_{d}$ and $\lambda_{s}$ in TDL. We set a fixed value of $\lambda_{s}$ at $0.05$ and vary $\lambda_{d}$. We find that using a larger $\lambda_{d}$ compared with $\lambda_{s}$ leads to better performance for DiChaViT. This suggests that knowing which channel a token comes from, i.e., the same or different channels, is necessary. The results indicate that imposing stricter constraints on tokens from different channels compared with tokens from the same channel obtains the best performance. Tab. 4 illustrates the impact of each component in TDL. We see that considering only tokens within the same channels (i.e., $\lambda_{d}=0$) is insufficient, resulting in a significant drop in performance, where using both components of DiChaViT yields the best performance.
Ablation on DCS. Tab. 5 shows the effect of temperature $t$ used in Algorithm 1 on DCS. When $t$ is set to a very small value, as reported in the first column ($\approx 0$), DCS selects channels with the lowest similarity scores to the anchor channel. Conversely, when $t$ is assigned a large value, denoted as HCS in the last column, DCS is reduced to HCS [18], meaning it selects the subset of channels randomly. We find that always selecting the lowest similar channels ($\approx 0$) does not yield the best performance. Instead, setting the temperature to $t=0.1$ produces favorable results for both So2Sat and CHAMMI.
DCS and HCS on the distribution of number sampling of the channels. Fig. 5 compares the number of times each channel is sampled during training with DCS (blue bars) and HCS [18] (red dashed line). DCS offers a different distribution for its channels compared to HCS, with some channels receiving more training than others. For example, Real part of Lee-filtered covariance matrix channel (last bar) is sampled twice as frequently as Band B8a channel (first bar).

### v2

@@ 3.3 Diverse Channel Sampling (DCS)
Bao et al. [18] introduced HCS to reduce the training time and improve the robustness of the model. The main concept is to randomly drop some input channels and train the model only on the remaining channels. In the same spirit, we propose a novel method, Diverse Channel Sampling (DCS), to sample a more diverse subset of channels during training (Fig. 2(c)). Similar to HCS, we start by randomly sampling a number $k$, which is the size of a subset of channels to train on. However, while HCS samples $k$ channels randomly, DCS first samples an anchor channel $c_{k}$. Then, we select other $k-1$ channels that are dissimilar to the anchor channel. This idea shares similarity with Channel DropBlock [53], where a set of similar channels in a CNN layer is masked out to disrupt co-adapted features. However, instead of keeping a fixed number of feature map channels as in Channel DropBlock, DCS selects a flexible number of input channels for each sampling. The procedure of DCS is outlined in Algorithm 1.
@@
Temperature $t_{\mathrm{DCS}}$
1 Sample a random variable $k$ uniformly from the set $\{1,2,...,m\}$
2 Sample an anchor channel $c_{k}$ uniformly from all $m$ channels
3 Compute the cosine similarity between channel $c_{k}$ and the other $m-1$ channels: $\mathbf{s}=[\left\langle f_{k},f_{i}\right\rangle,...]$, $\forall i\neq k$$(\mathbf{s}\in\mathbb{R}^{m-1})$
4 Convert $1-\mathbf{s}$ to probability using $\mathrm{softmax}$ with temperature $t_{\mathrm{DCS}}$:     $\mathbf{p}=\mathrm{softmax}((1-\mathbf{s})/t_{\mathrm{DCS}})$$(\mathbf{p}\in\mathbb{R}^{m-1})$
5 Sample $k-1$ distinct channels from $m-1$ channels with probability $\mathbf{p}$
6 Combine the $k-1$ channels with channel $c_{k}$ to create a set of $k$ sampled channels.
Output :  Image $X$ with only $k$ sampled channels
@@ 5 Sample $k-1$ distinct channels from $m-1$ channels with probability $\mathbf{p}$
In practice, Algorithm 1 can be applied to a batch of images for faster sampling. We use channel token $\mathrm{ch}_{i}$ to represent the channel feature $f_{i}$. Refer to Sec. 4.4 and Tab. 5 for more discussion on choices of $f$. The temperature $t_{\mathrm{DCS}}$ controls the sharpness of the probability distribution. With a large $t_{\mathrm{DCS}}$, DCS reduces to HCS, while with a small $t_{\mathrm{DCS}}$, DCS selects a random subset of channels that are the least similar to the anchor channel.
@@ 3.4 Training Objective
The final loss consists of the primary loss for the specific task (e.g., cross-entropy for classification), Channel Diversification Loss (CDL) applied to channel tokens, and Token Diversification Loss (TDL) used on patch tokens. These terms work together to promote diversity in channel and patch token features, resulting in a more robust model, as shown in Eq. 5:
@@ 4.1 Experimental Setup
  • DepthwiseViT [14] utilizes a depthwise convolution layer to independently filter each input channel. The resulting features are averaged to create a new feature representation, which is then fed into a ViT backbone.
@@
  • HyperNetViT [42] employs a neural network (e.g., MLP) to independently generate weights for each channel, which are then concatenated to form a patch projection layer. This patch projection layer is subsequently used in a ViT backbone.
  • ChAda-ViT [15] uses a shared projection layer to extract features from each channel separately, then feeds these tokens, together with their corresponding positional embeddings and channel embeddings, into a ViT backbone.
@@
Implementation details. As HCS proves robust in multi-channel imaging [18], we incorporate this technique for DepthwiseViT, TemplateMixingViT, and HyperNetViT to ensure a fair comparison in these adaptive baselines used by Chen et al. [14] [^1] https://github.com/chaudatascience/channel_adaptive_models. For ChannelViT and ChAda-ViT, due to their similarity (primarily a difference in whether HCS is included), we use the implementation from [18] for both methods [^2] https://github.com/insitro/ChannelViT. All baselines utilize a ViT small architecture ($21$M parameters) implemented in DINOv2 [54] as the backbone  [^3] https://github.com/facebookresearch/dinov2. We use AdamW optimizer [55] to train the models, minimizing cross-entropy loss on JUMP-CP and So2Sat, and proxy loss on CHAMMI. For the learning rate, we use a scheduler with linear warmup and cosine decay. Refer to Appendix Sec. A for details.
@@ 4.2 Datasets
So2Sat [17] contains synthetic aperture radar and multispectral optical image patches from remote sensing satellites. Each image in the dataset has $18$ channels, of which eight Sentinel-1 and $10$ Sentinel-2 channels. The dataset consists of $17$ classes, each representing a distinct climate zone. We use the city-split version of the dataset, which includes $352$K training images and $24$K test images.
@@ 4.3 Results
Table 1: Comparison of test accuracy of channel adaptive models. "Full" refers to inference on all channels, while "Partial" means testing on a subset of channels (Sentinel-1 channels for So2Sat, fluorescence channels for JUMP-CP). We find our model outperforms other baselines, with a $5.0$% boost on CHAMMI and a $1.5-2.5$% point improvement on JUMP-CP and So2Sat.
@@
DiChaViT (ours) 69.68 69.19 57.98 63.36 47.76
Table 2: Test accuracy of DiChaViT and ChannelViT on partial channels of JUMP-CP [12]. Each column represents mean±$\pm$±std for all combinations when tested on partial channels. For example, column "7" indicates testing on $7$ out of $8$ channels, and, thus, the reported variance is due to the presence or absence of a channel. See to Tab. 9 in the Appendix for detailed results for each combination for column "7" with model variance. DiChaViT consistently exhibits improved robustness in the presence of missing channels during inference.
@@
ChannelViT [18] 67.51 60.36$\pm$9.1 52.74$\pm$12.2 44.89$\pm$13.2 36.88$\pm$12.3 29.36$\pm$9.3 23.70$\pm$5.0 20.78$\pm$1.6
DiChaViT (ours) 69.19 61.91$\pm$9.3 54.49$\pm$12.4 46.35$\pm$13.4 38.00$\pm$12.4 30.09$\pm$9.3 23.97$\pm$4.9 20.90$\pm$1.6
Tab. 1 shows that DiChaViT outperforms the state-of-the-art ChannelViT by up to $5.0$% points on all three datasets: CHAMMI [14], JUMP-CP [12], and So2Sat [17]. For JUMP-CP and So2Sat, we consider two scenarios: tested on all training channels (denoted as "Full") and tested on a subset of channels (denoted as "Partial"). In the full channels setting, our model shows a $1.5-2.5$% point improvement compared with other baselines on JUMP-CP and So2Sat. When tested on partial channels, DiChaViT demonstrates its robustness by achieving a $1.5$% improvement compared with the baselines. This demonstrates that diversifying feature representations in MCI-ViT models boosts both performance and robustness.
Tab. 2 presents a detailed evaluation of DiChaViT and the best baseline model, ChannelViT, when tested on partial channels of the JUMP-CP dataset (with a total of eight channels). For the partial channel evaluation, we exclude some of the channels that the models were trained on and only test the model on the remaining channels. Then, we calculate the average accuracy across all combinations, e.g., testing on seven channels, as shown in column "7", involves averaging the results of $C^{7}_{8}=8$ combinations (refer to Tab. 9 in the Appendix for detailed results). Our findings consistently show that DiChaViT demonstrates improved robustness when some input channels are missing.
To provide more insight into the contribution of each component of DiChaViT, Tab. 3 presents the model’s performance when a component is removed. The results highlight the critical role of the DCS component, as its removal has the most detrimental effect on performance, particularly in the Partial setting, with a decrease of $16$% and $30$% points on JUMP-CP and So2Sat, respectively. The absence of CDL and TDL results in similar performance drops across all datasets. The highest scores are achieved when all components are integrated, indicating that each component plays a crucial role in the model’s design. Refer to Tab. 8 in the Appendix for a comprehensive analysis.
Table 3: Model ablations of DiChaViT. Removing any component in DiChaViT has a negative impact on overall performance, with significant decreases observed on the Partial setting when DCS is removed. Including all components improves performance across all three datasets.
@@
DiChaViT 69.66 69.19 57.98 63.36 47.76
w/o CDL 68.07 67.66 56.87 62.20 45.74
w/o TDL 67.61 68.12 56.62 62.39 46.87
w/o DCS 65.32 66.03 42.37 59.20 17.88
4.4 Analysis and Discussion
4.4.1 Role of Channel Tokens in MCI-ViT Models
The role of channel tokens. In MCI-ViT models such as ChannelViT [18] and ChAda-ViT [15], channel tokens play a crucial role in learning channel-specific features, particularly when dealing with multiple channels where each contains unique information. To assess the impact of channel tokens, we compared the performance of DiChaViT on JUMP-CP and CHAMMI with (orange bars) and without channel tokens (blue bars), as shown in Fig. 3. The results indicate that DiChaViT demonstrates significant improvements with channel tokens, resulting in $8.0$% and $15.0$% point increases on JUMP-CP and CHAMMI, respectively, highlighting their importance.
Figure 3: Performance of DiChaViT on JUMP-CP and CHAMMI with and without channel tokens. Using channel tokens with orthogonal initialization (green) improves performance.
Orthogonal initialization of channel tokens boosts performance. As shown in Fig. 3, using orthogonal initialization (green) provides a $1.0$% gain on JUMP-CP and CHAMMI. This may suggest that by initializing the weights orthogonally, the model can more effectively capture diverse patterns within the data, resulting in boosting its overall performance.
Table 4: Ablation on the two components of TDL. Only $\mathcal{L}_{\text{s}}$ indicates using only within channel tokens (i.e., $\lambda_{d}=0$), while Only $\mathcal{L}_{\text{d}}$ indicates the use of only tokens from different channels in Eq. 4. Incorporating both components in TDL gives the best performance.
@@ 4.4.1 Role of Channel Tokens in MCI-ViT Models
Both 63.36 69.66
Table 5: Different choices of channel feature $f$ in DCS (Algorithm 1). We compare the performance when using the channel tokens ($\mathrm{ch}_{i}$) and patch tokens (i.e., image patches after passing through the projection layer) to compute the similarity score for sampling.
So2Sat [17] CHAMMI [14]
Patch tokens 63.00 65.57
Channel tokens 63.36 69.68
Table 6: Effect of temperature $t_{\mathrm{DCS}}$ on DCS (Algorithm 1). The first column ($\approx 0$) indicates the use of a very small value of $t_{\mathrm{DCS}}$, which is reduced to selecting the lowest similarity channels. The last column indicates a large value of $t_{\mathrm{DCS}}$, which is reduced to HCS [18]. Using $t_{\mathrm{DCS}}=0.1$ obtain the best results on So2Sat and CHAMMI datasets.
Temperature $t_{\mathrm{DCS}}$ $\approx 0$ 0.001 0.01 0.1 0.2 HCS
@@
CHAMMI [14] 67.22 66.91 68.96 69.66 66.07 66.30
4.4.2 Ablation on Feature Diversification Losses (CDL and TDL)
Impact of $\lambda_{\mathrm{CDL}}$ (Eq. 5) in CDL. Fig. 4(a) and (b) show the performance of DiChaViT (mean and std) across different values of $\lambda_{\mathrm{CDL}}$ on So2Sat and CHAMMI datasets. We can observe that selecting a value that is too large is not beneficial to the performance. It is worth finding a suitable value for $\lambda_{\mathrm{CDL}}$. On the So2Sat, the best performance is achieved with $\lambda_{\mathrm{CDL}}=0.001$, while the suitable value for CHAMMI is $0.1$.
Ablation on TDL (Eq. 4). Fig. 4(c) reports the performance of our model across different ratios of $\lambda_{d}$ and $\lambda_{s}$ in TDL. We set a fixed value of $\lambda_{s}$ at $0.05$ and vary $\lambda_{d}$. We observe that using a larger $\lambda_{d}$ compared with $\lambda_{s}$ leads to better performance for DiChaViT. This suggests that knowing which channel a token comes from, i.e., the same or different channel, is necessary. The results indicate imposing stricter constraints on tokens from different channels compared with tokens from the same channel obtains the best performance. Tab. 4 shows the impact of each component in TDL. We see that considering only tokens within the same channels (denoted by "Only $\mathcal{L}_{\text{s}}$") is insufficient, resulting in a significant drop in performance. In contrast, using both $\mathcal{L}_{\text{s}}$ and $\mathcal{L}_{\text{d}}$ in TDL yields the best performance of DiChaViT.
4.4.3 Ablations for Diverse Channel Sampling (DCS)
Channel feature $f$ in DCS. Tab. 5 compares the performance of using channel tokens ($\mathrm{ch}_{i}$) and patch tokens (i.e., image patches after passing through the projection layer) to compute the similarity score for sampling in Algorithm 1 (line 3). We observe that using channel tokens gains better performance on So2Sat and CHAMMI datasets. Note that while channel tokens are shared across all input images, patch tokens differ for each input image.
Impact of temperature on DCS. Tab. 6 shows the effect of temperature $t_{\mathrm{DCS}}$ used in Algorithm 1 on DCS. When $t_{\mathrm{DCS}}$ is set to a very small value, as reported in the first column (denoted as "$\approx 0$"), DCS selects channels with the lowest similarity scores to the anchor channel. Conversely, when $t_{\mathrm{DCS}}$ is assigned a large value, denoted as "HCS" in the last column, DCS is reduced to HCS [18], meaning that it selects the subset of channels randomly. We find that always selecting the lowest similar channels ($\approx 0$) does not yield the best performance. Instead, setting the temperature to $t_{\mathrm{DCS}}=0.1$ produces favorable results for both So2Sat and CHAMMI.
DCS and HCS on the distribution of number sampling of the channels. Fig. 5 compares the number of times each channel is sampled during training with DCS (blue bars) and HCS [18] (red dashed line). DCS offers a different distribution for its channels compared with HCS, with some channels receiving more training than others. For example, Real part of Lee-filtered covariance matrix (Real Lee-Cov) in the last bar, is sampled twice as frequently as Band B8a channel (first bar).